OpenStudy (goformit100):

Find P(A/B) if P(A) = 0.4,P(B) = 0.8 andP(B/A) = 0.6

OpenStudy (goformit100):

@mikaela19900630 @noided

OpenStudy (anonymous):

Do you know formula for P(A/B) ??

OpenStudy (goformit100):

No Sir, How to do it ?

OpenStudy (anonymous):

Friend my Friend..

OpenStudy (anonymous):

I am not sure but it is I think: \[P(A/B) = \frac{A \cap B}{P(B)}\]

OpenStudy (goformit100):

ok

OpenStudy (nathan917):

Are you trying to find P(A/B)?

OpenStudy (goformit100):

ok I got how to do . Thanks

OpenStudy (nathan917):

If P(A) = 0.4 .P(B) =0.8 P(B/A) =0.6.Find P(A/B)? _____________________________________________________________________ By P(A), P(B), and P(B/A), we can find P(A^B), ie. A intersection B, coz P(B/A)=P(A^B)/P(A). And then substituting P(A^B) in P(A/B), ie. P(A/B)=P(A^B)/P(B). In your sum, Let, P(A^B)=x, Now, 0.6=x/0.4 So, x=0.24. So, P(A^B)=0.24. And, therefore, P(A/B)=0.24/0.8, ie. 0.3. Therefore, P(A/B)=0.3.

OpenStudy (nathan917):

Its kinda like this.

OpenStudy (anonymous):

\[P(A/B) = \frac{P(A \cap B)}{P(B)}\]

OpenStudy (goformit100):

Ok Thanks again

OpenStudy (anonymous):

\[P(B/A) = \frac{P(A \cap B)}{P(A)}\] Just divide them now..

OpenStudy (anonymous):

By taking ratio you get: \[\frac{P(A/B)}{P(B/A)} = \frac{P(A)}{P(B)}\] Now find what you want to find from this..

OpenStudy (goformit100):

ok Thanks @waterineyes